A student opens a book to two facing pages. The product of the page numbers is 1806, find the pages numbers.
First page__________
Second page_________
How do I work this problem?

Answer by Alan3354(69443)

A student opens a book to two facing pages. The product of the page numbers is 1806, find the pages numbers.
First page__________
Second page_________
How do I work this problem?
--------------------
The quickest and easiest way is to find the sqrt(1806)
= ~ 42.49
That's between the 2 numbers, so they're 42 and 43.
-----------------
The hard way:
n*(n+1) = 1806
n^2 + n - 1806 = 0
(n-42)*(n+43) = 0
n = 42